OpenStudy (anonymous):

Simplify –10 + (–3x) – 5 + 2x

OpenStudy (lilg132):

- 10 - 3x - 5 + 2x -15 - x

OpenStudy (anonymous):

Combine like terms. These would be the -10-5, making -15; and the -3x (do not forget the - sign if it was not in the parentheses) +2x, making -x so, -x-15 Put the highest power of your variable first in general

OpenStudy (anonymous):

i got -5 -5x

OpenStudy (anonymous):

Your signs were most likely mixed up. Be very careful of this

OpenStudy (anonymous):

-10-5-3x+2x was my set up matched the like terms

OpenStudy (lilg132):

- 10 - 5 = ?

OpenStudy (anonymous):

lindsay, you have it correct (-10-5-3x+2x) so now you say, "-10 - 5 = -15) (think of a numberline)

OpenStudy (anonymous):

–10 + (–3x) – 5 + 2x = -15-x